# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v1

@@ 5 Conclusion and Future Work
For future work, it would be interesting to investigate the application of deeper architectures for individual predictors and multimodal datasets. This would enable the construction of portfolios where the hypotheses are not limited to stock time-series but can include any source of information relevant to any particular hypothesis. Additionally, it would be valuable to test the proposed methodology in scenarios where the hypothesis is a group of stocks or a portfolio, and to explore how to choose an appropriate target in such cases. Another potential direction for future research is to examine the relationship between complexity and diversity in the ensemble loss within the PSEM framework. Finally, an interesting application especially in finance would be the creation of a risk premia factor that measures diversity by constructing different indices with varying levels of diversity and monitoring their risk-return profile across all diversity scales.

### v2

@@ 5 Conclusion and Future Work
For future work, it would be interesting to investigate the application of deeper architectures for individual predictors and multimodal datasets. This would enable the construction of portfolios where the hypotheses are not limited to stock time-series but can include any source of information relevant to any particular hypothesis. Additionally, it would be valuable to test the proposed methodology in scenarios where the hypothesis is a group of stocks or a portfolio, and to explore how to choose an appropriate target in such cases. Another potential direction for future research is to examine the relationship between complexity and diversity in the ensemble loss within the PSEM framework. Finally, an interesting financial application would be the development of a risk premia factor that takes long positions in high-diversity portfolios and short positions in low-diversity portfolios, aiming to profit from the diversity-quality trade-off in predicted returns presented in this work.